OpenStudy (anonymous):

8/9 = 96/? higher term

Parth (parthkohli):

HINTZ:\[\dfrac{8}{9} = \dfrac{8\cdot 12}{9 \cdot 12} = \dfrac{96}{\cdots}\]

OpenStudy (tyteen4a03):

For future questions like this, rewrite this relation in algebra form: \(\frac{8}{9} = \frac{96}{x}\). Now remember that \(\frac{a}{b} = \frac{c}{d}\) is equal to \(ad = cb\). (Multiply both sides by b then by d, and you will get this relation)